### v1

# Watson’s lemma revisited

Abstract.
In this paper, we study the operator analogue of Laplace’s method and the matrix form of Watson’s lemma. The main approach utilised is that described by Williams and Wong [J. Approx. Theory 24 (4) (1974), 378–384]. Some examples are also given.
Key words and phrases:
Asymptotics, Watson’s lemma, Laplace’s method, special matrix functions
2020 Mathematics Subject Classification:
33C47, 41A60, 44A10, 47A56
1. Introduction
Williams and Wong [9] studied the operator analogue of Watson’s lemma. In their work, a key assumption for the matrix form of Watson’s lemma is that the matrix is normal. Luke and Barker [5] tried to weaken the “normal” condition, but the constant $M_{1}$ in [5, Eq. (40)] is not absolute, which motivates us to find a correct answer.
This paper is organised as follows. In Section 2, we deduce an operator version of Laplace’s method. In Section 3, we make a revision of Luke and Barker’s result [5, Theorem 3]. This paper concludes with some examples in Section 4.
2. Operator analogue of Laplace’s method
In this section, we adopt the definition of asymptotic expansions of operator-valued functions introduced in [9, Sect. 3] and establish the operator analogue of Laplace’s method.
Let $X$ be a Banach space over the complex field $\mathbb{C}$ and $Y$ be a dense subspace of $X$. Let $A$ be a closed linear operator from $Y$ into $X$. The spectrum and resolvent of $A$ are denoted by $\sigma(A)$ and $R_{\lambda}(A)=\left(\lambda I-A\right)^{-1}$, respectively.
According to [9, Sect. 2], if, in addition, $A$ satisfies the following conditions:
  • $(\text{C}_{1})$ There exists a positive $\Delta$ such that
    $\sigma(A)\subset\left\{\lambda\in\mathbb{C}:\lambda\neq 0\text{ and }\left|\arg\lambda\right|\leqslant\frac{\pi}{2}-\Delta\right\},$
  • $(\text{C}_{2})$ Let $\omega(A):=\inf\{\Re(\lambda):\lambda\in\sigma(A)\}$. There exists $M>0$ and $0<\omega_{1}\leqslant\omega(A)$ such that for any positive integer $n$,
    $\left\|R_{\lambda}\left(A\right)^{n}\right\|\leqslant M\left(\omega_{1}-\lambda\right)^{-n}\quad(\lambda<\omega_{1}),$ (2.1)
then $e^{tA}$ is well-defined for real $t$. Then we can define for $0<\alpha<\infty$ that
$A^{-\alpha}=\frac{1}{\Gamma(\alpha)}\int_{0}^{\infty}t^{\alpha-1}e^{-tA}\mathrm{d}t.$ (2.2)
Furthermore, there exists a strongly continuous semigroup, $\left\{e^{-tA}:0\leqslant t<\infty\right\}$, of bounded linear operators on $X$ such that
$\left\|e^{-tA}\right\|\leqslant Me^{-t\omega_{1}},\quad\text{for all }0<t<\infty.$ (2.3)
Let us state our main results. Our proof is based on the classical procedure in [10, p. 57–60].
Theorem 2.1.
Let $\{A_{\alpha}\}$ be a net of closed linear operators from $Y$ into $X$, each satisfying conditions $(\text{C}_{1})$ and $(\text{C}_{2})$ with the same $\Delta$ and $M$ and such that there is some positive $\eta$ with $\omega_{1}(A_{\alpha})\geqslant\eta\,\omega(A_{\alpha})$ for each $A_{\alpha}$. Define the integral
$I(z):=\int_{a}^{b}\varphi(x)e^{-zh(x)}\mathrm{d}x,$
where $h(x)$ is real, $a\in\mathbb{R}$ is finite and $b(>a)$ is finite or infinite.
Assume that (i) $h(x)>h(a)$ for all $x\in(a,b)$, and for every $\delta>0$ the infimum of $h(x)-h(a)$ in $[a+\delta,b)$ is positive; (ii) $h^{\prime}(x)$ and $\varphi(x)$ are continuous in a neighborhood of $x=a$, except possibly at $a$; (iii) as $x\to a^{+}$, the following expansions hold
$h(x)\sim h(a)+\sum_{s=0}^{\infty}a_{s}\left(x-a\right)^{s+\mu},\quad\varphi(x)\sim\sum_{s=0}^{\infty}b_{s}\left(x-a\right)^{s+\lambda-1},$
where the expansion for $h(x)$ is differentiable, $\lambda$ and $\mu$ are positive constants, and $a_{0}>0$; and (iv) the integral $I(z)$ converges absolutely for all sufficiently large $z$.
Then the bounded linear operator $I(A_{\alpha})$ has the asymptotic expansion
$I(A_{\alpha})\sim e^{-h(a)A_{\alpha}}\sum_{s=0}^{\infty}\Gamma\left(\frac{s+\lambda}{\mu}\right)c_{s}A_{\alpha}^{-\frac{s+\lambda}{\mu}},\quad\text{as }\left\|A_{\alpha}^{-1}\right\|\to 0,$ (2.4)
where the coefficients $c_{s}$ are defined by the expansion
$\frac{\varphi(x)}{h^{\prime}(x)}\sim\sum_{s=0}^{\infty}c_{s}t^{(s+\lambda-\mu)/\mu},\quad\text{as }\,t=h(x)-h(a)\to 0^{+}.$ (2.5)
In particular, $c_{0}=b_{0}\mu^{-1}a_{0}^{-\lambda/\mu}$.
Proof.
Let $\omega_{1,\alpha}=\omega_{1}(A_{\alpha})$ and $\omega_{\alpha}=\omega(A_{\alpha})$. Then $\omega_{1,\alpha}^{-1}\leqslant\|A_{\alpha}^{-1}\|\left(\eta\sin\Delta\right)^{-1}$ (see [9, Eq. (3)]). Thus $\omega_{1,\alpha}\to+\infty$, as $\|A_{\alpha}^{-1}\|\to 0$. In view of this, using (2.3) and the condition (iv) gives
$\|I(A_{\alpha})\|$ $\leqslant\big{\|}e^{-h(a)A_{\alpha}}\big{\|}\int_{a}^{b}|\varphi(x)|\left\|e^{-(h(x)-h(a))A_{\alpha}}\right\|\mathrm{d}x$
$\leqslant M\big{\|}e^{-h(a)A_{\alpha}}\big{\|}\int_{a}^{b}|\varphi(x)|\cdot e^{-\omega_{1,\alpha}(h(x)-h(a))}\mathrm{d}x<\infty,$
which confirms the boundness of $I(A_{\alpha})$ for suitable $\alpha$.
Conditions (ii) and (iii) imply the existence of a number $c$ in $(a,b)$ such that $h^{\prime}(x)$ and $\varphi(x)$ are continuous in $(a,c]$, and $h^{\prime}(x)$ is positive there. Put $T=h(c)-h(a)>0$ and introduce the new variable $t=h(x)-h(a)$. Since $h(x)$ is increasing in $(a,c)$, we may write
$e^{h(a)A_{\alpha}}\int_{a}^{c}\varphi(x)e^{-h(x)A_{\alpha}}\mathrm{d}x=\int_{0}^{T}f(t)e^{-tA_{\alpha}}\mathrm{d}t$ (2.6)
with $f(t)$ beging the continuous function in $(0,T]$ given by
$f(t)=\varphi(x)\frac{\mathrm{d}x}{\mathrm{d}t}=\frac{\varphi(x)}{h^{\prime}(x)}.$
Now $f(t)$ has the expansion (2.5) as $t\to 0^{+}$. Applying the operator analogue of Watson’s lemma [9, Theorem 1] to the integral on the right-hand side of (2.6), we have
$\int_{a}^{c}\varphi(x)e^{-h(x)A_{\alpha}}\mathrm{d}x\sim e^{-h(a)A_{\alpha}}\sum_{s=0}^{\infty}\Gamma\left(\frac{s+\lambda}{\mu}\right)c_{s}A_{\alpha}^{-\frac{s+\lambda}{\mu}},\quad\text{as }\left\|A_{\alpha}^{-1}\right\|\to 0.$ (2.7)
Now estimate the integral on the remaining interval $(c,b)$. Let $\omega_{0}$ be a value of $z>0$ for which $I(z)$ is absolutely convergent, ant set
$\varepsilon=\inf_{c\leqslant x<b}(h(x)-h(a)).$
By condition (i), $\varepsilon$ is positive. Now for $\omega_{1,\alpha}\geqslant\omega_{0}$,
$\omega_{1,\alpha}[h(x)-h(a)]$ $=\left(\omega_{1,\alpha}-\omega_{0}\right)[h(x)-h(a)]+\omega_{0}[h(x)-h(a)]$
$\geqslant\left(\omega_{1,\alpha}-\omega_{0}\right)\varepsilon+\omega_{0}[h(x)-h(a)].$ (2.8)
Let $I_{1}(A_{\alpha}):=\int_{c}^{b}\varphi(x)e^{-h(x)A_{\alpha}}\mathrm{d}x$. Applying (2.3), (2.8) and the condition (iv), we get
$\big{\|}e^{h(a)A_{\alpha}}I_{1}(A_{\alpha})\big{\|}$ $\leqslant M\int_{c}^{b}|\varphi(x)|e^{-\omega_{1,\alpha}(h(x)-h(a))}\mathrm{d}x$
$\leqslant Me^{(\omega_{0}-\omega_{1,\alpha})\varepsilon}\int_{c}^{b}|\varphi(x)|e^{-\omega_{0}(h(x)-h(a))}\mathrm{d}x=Ke^{-\omega_{1,\alpha}\varepsilon}.$
Recall that $\omega_{1,\alpha}^{-1}\leqslant\|A_{\alpha}^{-1}\|\left(\eta\sin\Delta\right)^{-1}$. Then
$\big{\|}e^{h(a)A_{\alpha}}I_{1}(A_{\alpha})\big{\|}\leqslant Ke^{-C/\|A_{\alpha}^{-1}\|},$ (2.9)
where $C=\varepsilon\eta\sin\Delta$. The result (2.4) follows from (2.7) and (2.9). ∎
A similar discussion of [9, Corollary] gives the following result.
Corollary 2.2.
Let $\{A_{\alpha}\}$ be a net of bounded linear operators on a Hilbert space $\mathcal{H}$. If each $\{A_{\alpha}\}$ is normal and satisfies the condition $(\text{C}_{1})$ with the same $\Delta$, and if $\varphi(x)$ and $h(x)$ satisfy conditions (i), (ii), (iii) and (iv), then the expansion (2.4) holds.
3. The work of Luke and Barker revisited
Luke and Barker [5, Sect. II] showed that their result [5, Theorem 3] satisfies the hypotheses of [9, Theorem 1]. However, the constant $M_{1}$ in [5, Eq. (40)] depends on the index $m$. More explicitly, we have the following fact
$M_{1}=\sum_{r=0}^{p}\left(m\right)_{r}\left|\lambda_{1}-\lambda\right|^{-r}>1+m\left|\lambda_{1}-\lambda\right|^{-1},$
which implies that under their restrictions, the inequality (2.1) is not valid.
In this section, we shall make a revision of their result. Before this, let us do some preparations.
For $A\in\mathbb{C}^{m\times m}$, let $A^{+}:=\frac{1}{2}\left(A+A^{H}\right)$ and define its 2-norm by
$\|A\|=\sup_{\|x\|_{2}=1}\|Ax\|_{2},$
where for any vector $x\in\mathbb{C}^{m\times 1}$, $\|x\|_{2}=\sqrt{x^{H}x}$. Denote
$\mu(A)=\min\big{\{}z\in\mathbb{C}:z\in\sigma\left(A^{+}\right)\big{\}},\ \omega(A)=\min\{\Re(\lambda):\lambda\in\sigma(A)\}.$
By [3, p. 647], it follows that
$\left\|e^{-tA}\right\|\leqslant e^{-t\mu(A)}\quad(t\geqslant 0).$ (3.1)
Next we derive the matrix form of Watson’s lemma.
Theorem 3.1.
Let $A$ be a matrix such that the condition $(\text{C}_{1})$ (see Sect. 2) is satisfied and there is a positive $\eta$ with $\mu(A)\geqslant\eta\,\omega(A)$. Assume that: (a) the function $f(t)$ is a locally integrable function on $[0,\infty)$ and admits the expansion
$f(t)=\sum_{n=1}^{\infty}a_{n}t^{n/r-1},\quad|t|\leqslant c+\delta,$
where $r,c$ and $\delta$ are positive; and (b) there exist positive constants $M_{0}$ and $b$ independent of $t$ such that $|f(t)|<M_{0}e^{bt}$ holds for $t\geqslant c$. Then
$\int_{0}^{\infty}f(t)e^{-tA}\mathrm{d}t\sim\sum_{n=1}^{\infty}a_{n}\Gamma(n/r)A^{-n/r},\quad\text{as }\left\|A^{-1}\right\|\to 0.$ (3.2)
Proof.
Fix an integer $N\geqslant 2$. Clearly, there is a constant $C$ such that for all $t>0$,
$\bigg{|}f(t)-\sum_{n=1}^{N-1}a_{n}t^{n/r-1}\bigg{|}\leqslant Ct^{N/r-1}e^{bt}.$ (3.3)
A similar derivation of [9, Eq. (3)] gives
$\mu(A)\geqslant\left\|A^{-1}\right\|^{-1}\eta\sin\Delta.$ (3.4)
This together with (3.1) and (2.2) shows that $A^{-n/r}$ is well-defined. Using (2.2), we can infer that
$R_{N}(A):=\int_{0}^{\infty}f(t)e^{-tA}\mathrm{d}t-\sum_{n=1}^{N-1}a_{n}\Gamma(n/r)A^{-n/r}$
satisfies
$R_{N}(A)=\int_{0}^{\infty}\bigg{(}f(t)-\sum_{n=1}^{N-1}a_{n}t^{n/r-1}\bigg{)}e^{-tA}\mathrm{d}t.$
If $\|A^{-1}\|\leqslant\frac{\eta\sin\Delta}{2b}$, then $\mu(A)\geqslant 2b$ by (3.4) and hence
$\left\|R_{N}(A)\right\|\leqslant CM\int_{0}^{\infty}t^{N/r-1}e^{(b-\mu(A))t}\mathrm{d}t$
by virtue of (3.1) and (3.3). Since $\mu(A)\to\infty$, it follows from (3.4) that
$\left\|R_{N}(A)\right\|=\mathcal{O}\left(\mu\left(A\right)^{-N/r}\right)=\mathcal{O}\left(\left\|A^{-1}\right\|^{-N/r}\right).$
Recall the inequality [9, p. 382]
$\left\|A^{-1}\right\|^{-N/r}\leqslant C_{1}\big{\|}A^{-(N-1)/r}\big{\|}\left\|A^{-1}\right\|^{1/r},$
where $C_{1}=C_{1}(N,r)$. Thus $\left\|R_{N}(A)\right\|=o\left(\left\|A^{-(N-1)/r}\right\|\right)$, which completes the proof. ∎
Corollary 3.2.
Let $A$ be a normal matrix satisfying the condition $(\text{C}_{1})$ and $f(t)$ be a function satisfying the conditions (a) and (b). Then the expansion (3.2) holds.
Proof.
It suffices to show that $\mu(A)\geqslant\eta\,\omega(A)$. Indeed, the further fact that $\mu(A)=\omega(A)$ follows from the spectral theorem for normal matrices. ∎
4. Examples
In this section, we deal with asymptotics of the gamma matrix function and Bessel matrix functions. Here we assume that $A$ is a matrix in $\mathbb{C}^{m\times m}$ and the norm $\|A\|$ is a least upper bound on the matrices of complex numbers.
4.1. Gamma matrix function
If $A$ is a matrix such that $\Re(z)>0$ for any eigenvalue $z$ of $A$, then the gamma matrix function $\Gamma(A)$ is defined as [4, Eq. (1)]
$\Gamma(A)=\int_{0}^{\infty}t^{A-I}e^{-t}\mathrm{d}t,\quad t^{A-I}\equiv\exp\left((A-I)\log t\right).$
Moreover, if $A+nI$ is invertible for all integer $n\geqslant 0$, then $\Gamma(A)$ is invertible, and the reciprocal gamma function is defined as
$\Gamma^{-1}(A)=A(A+I)\cdots(A+(n-1)I)\Gamma\left(A+nI\right)^{-1},\quad n\geqslant 1.$
Clearly, $\Gamma^{-1}(A)=\Gamma\left(A\right)^{-1}$ if $\sigma(A)\subset\mathbb{C}\backslash\mathbb{Z}_{\leqslant 0}$.
Now we can establish asymptotics of the gamma matrix function.
Theorem 4.1.
Let $A$ be a matrix satisfying: (c) $\sigma(A)\subset\{\lambda\in\mathbb{C}:\lambda\neq 0,\left|\arg(\lambda-\rho)\right|<\Delta\}$ with $0<\Delta<\frac{\pi}{2}$ and $0<\theta<2\pi$; and (d) for an invertible matrix $P$ such that $P^{-1}AP=J$ is the Jordan normal form of $A$, the number $n(P):=\|P\|\|P^{-1}\|$ is bounded as $\|A^{-1}\|\to 0$. Then
$\Gamma(A)\sim\sqrt{2\pi}\,e^{-A}A^{A-I/2}\sum_{k=0}^{\infty}g_{k}A^{-k},\quad\text{as }\left\|A^{-1}\right\|\to 0,$ (4.1)
where $g_{k}$ has explicit expressions. In particular, $g_{0}=1$ and $g_{1}=\frac{1}{12}$.
Proof.
Note that the scaled gamma function
$\Gamma^{*}(z):=\frac{1}{\sqrt{2\pi}}e^{z}z^{1/2-z}\Gamma(z)$
is analytic in the domain described in the condition (c). Since (4.1) holds for $\Gamma^{*}(z)$ as $|z|\to\infty$, we can make use of [5, Theorem 1] and obtain that (4.1) holds for $\Gamma^{*}(A)$ as $\|A^{-1}\|\to 0$. ∎
Theorem 4.2.
If $A$ is Hermitian and postive definite, then for any integer $N\geqslant 1$,
$\Gamma(A)=\sqrt{2\pi}\,e^{-A}A^{A-I/2}\bigg{(}\sum_{k=0}^{N-1}g_{k}A^{-k}+R_{N}(A)\bigg{)},$
and the reminder $R_{N}(A)$ has the estimate
$\left\|R_{N}(A)\right\|_{2}\leqslant\frac{1+\zeta(N)}{\left(2\pi\right)^{N+1}}\Gamma(N)\left\|A^{-1}\right\|_{2}^{-N},$ (4.2)
where $\zeta$ denotes the Riemann zeta function.
Proof.
Now take the spectral norm for matrices: $\|A\|\equiv\|A\|_{2}$. According to [5, p. 100] (see also [7, p. 271]), it holds for $\|B\|_{2}$ that if $B$ is Hermitian and positive definite, then any inequality satisfied by $f(z)$ must also be satisfied by $f(B)$. Thus, (4.2) follows from the error bound $R_{N}(z)$ for asymptotics of the scalar gamma function(see [1]) and the fact that $A$ is positive definite. ∎
Remark 4.1.
  • (1)
    Karamata’s Tauberian theorem is a converse of Watson’s lemma (see [2, Sect. XIII.5]) and its matrix analogue was given in Meerschaert and Scheffler [6].
  • (2)
    It is of interest to find the matrix analogues of asymptotic methods for contour integrals, since these methods may be helpful to derive asymptotics of special matrix functions, such as the asymptotics of $\Gamma(A)$ in the general case $\sigma(A)\subset\{\lambda\in\mathbb{C}:\lambda\neq 0,\left|\arg\lambda\right|<\pi-\delta\}$.
4.2. Bessel matrix functions
Integral expressions of Bessel matrix functions are given by
$J_{A}(z)$ $=\left(\frac{z}{2}\right)^{A}\frac{\Gamma^{-1}\left(A+I/2\right)}{\sqrt{\pi}}\int_{-1}^{1}\left(1-t^{2}\right)^{A-I/2}\cos(zt)\mathrm{d}t,\ \ \left|\arg z\right|<\pi;$
$I_{A}(z)$ $=\left(\frac{z}{2}\right)^{A}\frac{\Gamma^{-1}\left(A+I/2\right)}{\sqrt{\pi}}\int_{-1}^{1}\left(1-t^{2}\right)^{A-I/2}\cosh(zt)\mathrm{d}t,\ \ \left|\arg z\right|<\frac{\pi}{2},$
where $z\neq 0$ and $\mu(A)>-\frac{1}{2}$; see [8, Sect. 3] for details.
Let $A$ be a normal matrix satisfying the condition $(C_{1})$ and set
$h(t)=-\log\left(1-t^{2}\right),\ L(A)=\left(\frac{z}{2}\right)^{A}\Gamma^{-1}\left(A+I/2\right)\left(A-I/2\right)^{-1/2}.$
Then Theorem 2.1 confirms that as $\|A^{-1}\|\to 0$,
$J_{A}(z)\sim L(A),\ z\neq 0,\left|\arg z\right|<\pi;\quad I_{A}(z)\sim L(A),\ z\neq 0,\left|\arg z\right|<\frac{\pi}{2}.$ (4.3)
Since $A$ is normal, the fact that $n(P)=1$ verifies for the spectral norm and thus $n(P)$ is bounded for any matrix form. Hence, applying Theorem 2.1 and (4.3), we get that as $\|A^{-1}\|\to 0$,
$J_{A}(z)$ $\sim\frac{1}{\sqrt{2\pi}}\left(\frac{ez}{2}\right)^{A}A^{-A-I/2},\quad z\neq 0,\left|\arg z\right|<\pi;$
$I_{A}(z)$ $\sim\frac{1}{\sqrt{2\pi}}\left(\frac{ez}{2}\right)^{A}A^{-A-I/2},\quad z\neq 0,\left|\arg z\right|<\frac{\pi}{2}.$